OpenStudy (anonymous):

What is the equation of the line passing through (2, –2) and perpendicular to the line y = –x + 3 in slope-intercept form?

OpenStudy (anonymous):

y=x-4

OpenStudy (anonymous):

In slope intercept form Y=x-4 :)

OpenStudy (anonymous):

First find the slope of the desired line. Since the given line has a slope of -1, the perpendicular to that will have a slope of 1 (the opposite of the reciprocal). Then put the slope and point into the point-slope form, and simplify to slope-intercept form, thus\[y+2=1(x-2) \implies y=x-4\]

OpenStudy (anonymous):

Do math every day.